OpenStudy (hitaro9):

Suppose 2 is a primitive root of 29. Find all solutions to 4x^12 = -23 mod 29

OpenStudy (hitaro9):

I've reduced it down (barely) to 2x^12 = 3 mod 29 But I'm not sure how to proceed without basically recreating an entire indice table.

OpenStudy (dan815):

can you explain primitive root

OpenStudy (hitaro9):

There is some k such that 2^k is congruent to j for all j relatively prime to 29.

OpenStudy (rational):

take discrete logarithm both sides and isolate x

OpenStudy (hitaro9):

So if I have like ind2(3) mod 28 How would I even solve that?

OpenStudy (rational):

\[2x^{12}\equiv 3\pmod{29}~~~ \iff~~~ \text{ind}_2(2x^{12})\equiv \text{ind}_2(3) \pmod{28}\]

OpenStudy (hitaro9):

I'd have 2^k = 3 mod 28 for some k, but I wouldn't know how to find that k without guessing and checking/ without a table?

OpenStudy (rational):

you don't need a table to compute \(\text{ind}_2(2)\) because \(\text{ind}_b b=1\)

OpenStudy (hitaro9):

Right that makes sense

OpenStudy (rational):

use logarithm properties and break it down as much as possible first

OpenStudy (hitaro9):

I'm not following. What logarithm properties should I be considering?

OpenStudy (rational):

\[\text{ind}(mn) = \text{ind}(m) + \text{ind}(n)\] and \[\text{ind}(m^n) = n*\text{ind}(m) \]

OpenStudy (hitaro9):

Right so I'd have something like 1+12ind2(x) = ind2(3) mod 28 I don't know how to deal with the right side though

OpenStudy (rational):

looks good, next you need to find \(\text{ind}_2(3)\)

OpenStudy (rational):

you need to work it, there is no other way : 2^2 = 4 2^3 = 8 2^4 = 16 2^5 = 32 = 3 so the index of 3 is 5

OpenStudy (rational):

i mean to say there is no other way, you need to try the exponents till you get what you want

OpenStudy (hitaro9):

Oh okay then. Fair enough

OpenStudy (rational):

rest should be easy

OpenStudy (rational):

1+12ind2(x) = ind2(3) mod 28 1+12ind2(x) = 5 mod 28 12ind2(x) =4 mod 28 3ind2(x) =1 mod 7 ind2(x) = 5 mod 7

OpenStudy (rational):

so we get \[\text{ind}_2(x) = 5, 12, 19, 26\]

OpenStudy (rational):

consequently the solutions would be \[x \equiv 2^5,~2^{12},~2^{19},~2^{26}\pmod{29}\]

OpenStudy (hitaro9):

Right, that all makes sense. Thank you so much for everything you've done today. You've been absolutely phenomenal

OpenStudy (rational):

yeah that keeps me going for trying ur next problem ;p yw :)

OpenStudy (hitaro9):

Haha gotta sweet talk ya' to make sure you do all the problems. You've figured me out. But for real though, you've been amazing thank you.